Give the equation of a conic which is symmetric to the line x+2 = 0.
1
Expert's answer
2020-02-19T11:02:00-0500

The conic symmetric to line x+2=0 would be a parabola with the given line acting as its axis.

Let the vertex of this parabola lie on the x-axis itself. Hence, it is of the form "(a,0)". Clearly, the vertex also lies on the axis as well.

"\implies a=-2 \implies (-2,0)" is the vertex of the parabola.

Equation of a parabola with axis x=0 is "y=x^2". Thus, by shifting origin to (-2,0), we get

"y=(x+2)^2"

"\implies y=x^2+4x+4" is the required equation of the conic which is symmetric to x+2=0